# Area Selective Chemical Vapor Deposition of Gold by Electron Beam Seeding Zusammenfassung

### v1

@@ Conclusions
In this work, area selective CVD of gold by electron beam seeding is demonstrated for the first time. Gold nuclei, acting as a seed layer, are deposited through a maskless procedure by means of a focused electron beam, followed by gentle heating of the substrate, which triggers autocatalytic gold growth only in the seeded regions. The use of a focused electron beam ensures unparalleled spatial resolution, unhampered by mask usage typical in electron beam lithography and rendering complex multi-stage lithographic processing unnecessary. Seeding and subsequent growth both use the same precursor compound, Au(acac)Me, which is commercially available, long-term stable and non-toxic. The technique has been demonstrated to operate at low temperatures of approximately 80 °C, more than 200 K below the typical substrate temperature used in the established CVD process for this precursor. Moreover, it demands solely a high vacuum, making technological implementation remarkably straightforward. Any commercially available scanning electron microscope can be used for this method if it is equipped with a gas injection system and a heating stage. In the future, this process could be seamlessly integrated into conventional CVD reactors by equipping them with an electron gun featuring programmable position control.
@@
Unlike many other area selective CVD processes, direct electron beam writing for seeding is not affected by the surface material or topography. To demonstrate this capability, the same precursor was used to fabricate vertical nanostructures with high aspect ratios using the focused electron beam. In addition to achieving autocatalytically active seeds in 3D, the first quantitative measurement of the temperature change during direct electron beam writing of 3D structures was performed. This quantitative determination of the up to now inaccessible local temperature evolution is a major breakthrough as it enables improved modelling for accurate shape 3D nanoprinting through FEBID within the additive manufacturing field at a sub-100 nm scale 47, 49, 44, 45, 46.
@@
To summarise, this work presents a new, technologically powerful and easy to implement approach to area selective gold deposition. The low temperature necessary makes it highly suitable for processing innovative flexible devices that rely on heat-sensitive polymers, including flexible electronics 50 and flexible solar cells 51. Another exciting area of application is utilizing the maskless direct write seeding on 3D architectures, which enables the fabrication of nanostructures or localized electric contacts on intricate 3D devices.
@@ Acknowledgements
This work was funded by the German Research Association under grant agreement no. HO 5461/3-1 and by the Swiss National Fund by the COST-SNF project IZCOZ0_205450. We want to furthermore acknowledge support by the EU COST action CA 19140 ’FIT4NANO’ (www-fit4nano.eu). The authors are particularly grateful to Caroline Hain for her support with the English language.
Conflict of Interest
Authors declare no conflict of interest.
Keywords:
Literatur
- [1] T. H. Baum, P. B. Comita, Chemical Vapor Deposition of Gold and Silver, chapter 6, pages 303–327, John Wiley & Sons, Ltd 1994.
- [2] T. T. Kodas, M. J. Hampden-Smith, Overview of Metal CVD, chapter 9, pages 429–498, John Wiley & Sons, Ltd 1994.
- [3] S. O. Colgate, G. J. Palenik, V. E. House, D. W. Schoenfeld, Journal of Vacuum Science & Technology A 1990, 8, 1411.
- [4] M. M. Banaszak Holl, P. F. Seidler, S. P. Kowalczyk, F. R. McFeely, Applied Physics Letters 1993, 62, 1475.
- [5] B. Lecohier, B. Calpini, J. Philippoz, H. van den Bergh, Journal of Applied Physics 1992, 72, 2022.
- [6] W. L. Gladfelter, Chemistry of Materials 1993, 5, 1372.
- [7] T. H. Baum, C. R. Jones, Applied Physics Letters 1985, 47, 538.
- [8] S. J. C. Irvine, D. Lamb, Photo-assisted CVD, in Chemical Vapour Deposition: Precursors, Processes and Applications, The Royal Society of Chemistry 2008.
- [9] J. Lasseter, P. D. Rack, S. J. Randolph, Nanomaterials 2023, 13.
- [10] F. Porrati, S. Barth, G. C. Gazzadi, S. Frabboni, O. M. Volkov, D. Makarov, M. Huth, ACS Nano 2023, 17, 4704, pMID: 36826847.
- [11] R. R. Kunz, T. E. Allen, T. M. Mayer, Journal of Vacuum Science & Technology B: Microelectronics Processing and Phenomena 1987, 5, 1427.
- [12] R. R. Kunz, T. M. Mayer, Applied Physics Letters 1987, 50, 962.
- [13] M.-M. Walz, M. Schirmer, F. Vollnhals, T. Lukasczyk, H.-P. Steinrück, H. Marbach, Angewandte Chemie International Edition 2010, 49, 4669.
- [14] K. Muthukumar, H. O. Jeschke, R. Valentí, E. Begun, J. Schwenk, F. Porrati, M. Huth, Beilstein Journal of Nanotechnology 2012, 3, 546.
- [15] R. R. Kunz, T. M. Mayer, Journal of Vacuum Science & Technology B: Microelectronics Processing and Phenomena 1988, 6, 1557.
- [16] A. J. M. Mackus, J. J. L. Mulders, M. C. M. van de Sanden, W. M. M. Kessels, Journal of Applied Physics 2010, 107, 116102.
- [17] A. Mameli, B. Karasulu, M. A. Verheijen, B. Barcones, B. Macco, A. J. M. Mackus, W. M. M. E. Kessels, F. Roozeboom, Chemistry of Materials 2019, 31, 1250.
- [18] I. Utke, P. Hoffmann, J. Melngailis, J. Vac. Sci. Technol. B 2008, 26, 1197.
- [19] W. F. van Dorp, C. W. Hagen, P. A. Crozier, P. Kruit, Nanotechnology 2008, 19, 225305.
- [20] F. Vollnhals, M. Drost, F. Tu, E. Carrasco, A. Späth, R. H. Fink, H.-P. Steinrück, H. Marbach, Beilstein Journal of Nanotechnology 2014, 5, 1175.
- [21] P. Martinović, L. Barnewitz, M. Rohdenburg, P. Swiderek, Journal of Vacuum Science & Technology A 2023, 41, 033207.
- [22] E. Feurer, H. Suhr, Applied Physics A 1987, 44, 171.
- [23] C. E. Larson, T. H. Baum, R. L. Jackson, MRS Online Proceedings Library (OPL) 1986, 75, 721.
- [24] A. Botman, J. J. L. Mulders, C. W. Hagen, Nanotechnology 2009, 20, 17pp.
- [25] S. Barth, M. Huth, F. Jungwirth, J. Mater. Chem. C 2020, 8, 15884.
- [26] J. J. L. Mulders, J. M. Veerhoek, E. G. T. Bosch, P. H. F. Trompenaars, Journal of Physics D: Applied Physics 2012, 45, 475301.
- [27] I. Utke, P. Hoffmann, B. Dwir, K. Leifer, E. Kapon, P. Doppelt, Journal of Vacuum Science & Technology B: Microelectronics and Nanometer Structures 2000, 18, 3168.
- [28] A. Marashdeh, T. Tiesma, N. J. C. van Velzen, S. Harder, R. W. A. Havenith, J. T. M. De Hosson, W. F. van Dorp, Beilstein Journal of Nanotechnology 2017, 8, 2753.
- [29] D. McIntosh, G. A. Ozin, Inorganic Chemistry 1977, 16, 51.
- [30] L. Jiang, Q. Xu, The Journal of Physical Chemistry A 2005, 109, 1026.
- [31] I. Utke, P. Swiderek, K. Höflich, K. Madajska, J. Jurczyk, P. Martinović, I. Szymańska, Coordination Chemistry Reviews 2022, 458, 213851.
- [32] R. Winkler, J. D. Fowlkes, P. D. Rack, H. Plank, Journal of Applied Physics 2019, 125, 210901.
- [33] O. Sqalli, I. Utke, P. Hoffmann, F. Marquis-Weible, Journal of Applied Physics 2002, 92, 1078.
- [34] J. J. L. Mulders, L. M. Belova, A. Riazanova, Nanotechnology 2010, 22, 055302.
- [35] K. Höflich, R. B. Yang, A. Berger, G. Leuchs, S. Christiansen, Adv. Mater. 2011, 23, 2657.
- [36] R. Winkler, F. P. Schmidt, U. Haselmann, J. D. Fowlkes, B. B. Lewis, G. Kothleitner, P. D. Rack, H. Plank, ACS Applied Materials and Interfaces 2017, 9, 8233.
- [37] D. Kuhness, A. Gruber, R. Winkler, J. Sattelkow, H. Fitzek, I. Letofsky-Papst, G. Kothleitner, H. Plank, ACS Applied Materials and Interfaces 2021, 13, 1178.
- [38] K. Höflich, J. Jurczyk, M. Puydinger, Y. Zhang, C. Kapusta, I. Utke, M. Puydinger dos Santos, M. Götz, C. Guerra-Nuñez, J. Best, C. Kapusta, I. Utke, ACS Appl. Mater. Interfaces 2017, 9, 24071.
- [39] P. Martinović, M. Rohdenburg, A. Butrymowicz, S. Sarigül, P. Huth, R. Denecke, I. B. Szymanska, P. Swiderek, Nanomaterials 2022.
- [40] G. Hochleitner, H. D. Wanzenboeck, E. Bertagnolli, Journal of Vacuum Science & Technology B: Microelectronics and Nanometer Structures Processing, Measurement, and Phenomena 2008, 26, 939.
- [41] R. Córdoba, J. Sesé, M. Ibarra, J. De Teresa, Applied Surface Science 2012, 263, 242.
- [42] M. Haruta, Catal. Today 1997, 36, 153 .
- [43] S. J. Randolph, J. D. Fowlkes, P. D. Rack, Journal of Applied Physics 2005, 97, 124312.
- [44] E. Mutunga, R. Winkler, P. D. Rack, H. Plank, J. D. Fowlkes, ACS Nano 2019, 13, 5198 .
- [45] L. Skoric, D. Sanz-Hernández, F. Meng, C. Donnelly, S. Merino-Aceituno, A. Fernández-Pacheco, Nano Letters 2020, 20, 184.
- [46] J. D. Fowlkes, R. Winkler, P. D. Rack, H. Plank, ACS Omega 2023, 8, 3148 .
- [47] J. D. Fowlkes, R. Winkler, B. B. Lewis, M. G. Stanford, H. Plank, P. D. Rack, ACS Nano 2016, 10, 6163.
- [48] J. Robertson, Mater. Sci. Eng. R 2002, 37, 129.
- [49] L. Keller, M. Huth, Beilstein Journal of Nanotechnology 2018, 9, 2581.
- [50] B. D. Gates, Science 2009, 323, 1566.
- [51] M. Pagliaro, R. Ciriminna, G. Palmisano, Chem. Sus. Chem 2008, 1, 880 .
- [52] K. Rykaczewski, W. B. White, A. G. Fedorov, Journal of Applied Physics 2007, 101, 054307.
- [53] D. C. Joy, Monte Carlo modeling for electron microscopy and microanalysis, volume 9, Oxford University Press 1995.
- [54] R. Browning, T. Z. Li, B. Chui, J. Ye, R. F. W. Pease, Z. Czyżewski, D. C. Joy, J. Appl. Phys. 1994, 76, 2016.
- [55] L. Reimer, Measurement Science and Technology 2000, 11, 1826.
@@
Methods and Full Seed Array
Silicon wafers with native oxide and varied doping levels were employed as the substrate material and chipped into 1 cm x 1 cm pieces. As the doping level decides the conductivity and therefore the secondary electron yields, the minimum required electron doses for successful seeding varied. The silicon chips with a native oxide layer were vertically sonicated in N-Methyl-2-pyrrolidone for 20 min, followed by a rinse with isopropyl alcohol. The chips were then soaked in a 3:1 Piranha solution for at least 1 h before being rinsed with deionized water and dried using nitrogen gas. This process effectively removes any organic material from the surface of the sample.
All testing was carried out using a Helios 5 UX dual beam microscope. The gas injection system (GIS) with crystals of Au(acac)Me (CAS: 14951-50-9) was heated to 30 °C. To enable autocatalysis growth of gold films, gold nuclei were seeded with an electron beam at 5 kV acceleration voltage and a beam current of 100 pA. The base chamber pressure prior to opening the GIS valve was typically mbar, with the GIS valve open during autocatalysis - mbar at the start of the experiments, dropping to approximately at the end of an experimental day. The Kleindiek vacuum compatible Micro Heating and Cooling Stage (MHCS) was used together with a temperature control system to heat the sample up to 110 °C. The complete pattern array for all experiments is shown in Figure SI.1(a). It comprises 2 µm x 2 µm squares with variable pixel-to-pixel pitches (5, 10 and 20 nm) of the electron beam, separated by 2 µm, and different deposition times (achieved by scan repeats from 1 to 100). A non-linear color scale indicates the deposition time of a particular square, ranging from 10 ms to 16 s. The whole seed array takes 1 min 18 s of deposition time to complete. The assortment of squares within the full seed array enables the gold film thickness to be optimised and to understand how autocatalytic growth depends on the electron beam patterning. Squares featuring the same deposition time (and dose) are highlighted with a coloured frame.
The experimental layout of all autocatalytically grown films is shown schematically in Figure SI.1(b). The complete experimental procedure for seeding and autocatalytic growth at various temperatures is detailed below:
@@
Move the stage to the first patterning area, which should be at least 50 µm away from the focusing area (position i in Figure SI.1(b)), without conducting any observation with the SEM.
@@
Moving to the second patterning area, which is situated 100 µm from the previously grown films (position ii in Figure SI.1(b)).
@@
Pillars for measuring the temperature induced by an electron beam were grown using a focused electron beam of 15 kV and 100 pA. Deposition time for all pillars was achieved by a spot deposition with 5725 repeats with the maximum possible dwell time of 26.2 ms, resulting in a total deposition time of about 2 min 30 s for a single pillar. The layout for experimental deposition of all pillars under different stage temperatures is illustrated in Figure SI.1(c). The entire experimental protocol for pillars growth is described below:
@@
Without observation with the SEM, move the stage to the first patterning area, which is at least 20 µm away from the focusing area. Room temperature deposition is done at position i in Figure SI.1(c).
@@
Repeat the same protocol for higher temperatures ranging from 40 to 110 °C with a step of 10 K.
@@
SE Micrographs of Autocatalytically Grown Films on the Full Seed Arrays
@@
SE micrographs of gold films, grown on the full seed arrays at various temperatures, are displayed in Figure SI.2. The gold films in the top row were grown on seed arrays deposited at a room temperature of 30 °C, while the seed arrays for the gold films in the bottom row were deposited at higher temperatures ranging from 60 to 90 °C. Autocatalytic growth was continued for 30 min for each film.
Time Evolution of Autocatalytic Growth on the Full Seed Arrays
@@
The autocatalytic growth of gold films on two seed arrays for additional intermediate electron supply is shown in Figure SI.3, indicating the time evolution. The initial SEM image at 0 minutes in the top left corner shows the as-deposited seed array at room temperature (30 °C), taken immediately after heating to 90 °C. With the GIS open and stage heated for 10 min, several images of this array were captured every 60 seconds. The images were taken with a field of view of 40 µm x 26.6 µm for a total time of 23.9 s. Thus, 100 pA electron beam current resulted in a dose of 2.2 pC/µm or 14 electrons per nm per single image. Figure SI.3 displays SEM images captured at 3, 7, and 10 mins to visualize the evolution of the autocatalytic gold CVD film growth. The bottom row presents the as-deposited seed array, which was grown on the sample at an elevated temperature of 90 °C, with the same sequence of images taken at different times. These images show that SE imaging acts as iterative seeding to enhance the CVD growth rate significantly compared to the pure autocatalytic growth demonstrated above. Note that when imaging, seeds are also deposited between squares, resulting in a decrease in area-selectivity. By implementing an iterative seeding approach, electrons will only be supplied in the selected areas.
Autocatalytic Control Test
A control test of pattern array deposition was performed without the Au(acac)Me precursor gas being introduced into the vacuum chamber. The array pattern was irradiated with the electron beam while residual chamber gas was present at a base pressure of mbar, both at room temperature (30 °C) and at 90 °C. Subsequently, both irradiated areas were left at elevated temperature of 90 °C for 50 min with the gold precursor gas flow enabled. SE micrographs presenting the results are displayed in Figure SI.4(a) and (b).

### v2

@@ 3 Conclusions and Outlook
In this work, area selective CVD of gold by electron beam seeding is demonstrated for the first time. Gold nuclei, acting as a seed layer, are deposited through a maskless procedure by means of a focused electron beam, followed by gentle heating of the substrate, which triggers autocatalytic gold growth only in the seeded regions. The use of a focused electron beam ensures high spatial resolution, unhampered by mask usage typical in electron beam lithography and rendering complex multi-stage lithographic processing unnecessary. Seeding and subsequent growth both use the same precursor compound, Au(acac)Me2, which is commercially available, long-term stable and non-toxic. The technique has been demonstrated to operate at low temperatures of approximately 80 °C, more than 200 K below the typical substrate temperature used in the established CVD process for this precursor. Moreover, it demands solely a high vacuum, making technological implementation remarkably straightforward for academic research prototyping. Any commercially available scanning electron microscope can be used for this method if it is equipped with a gas injection system and a heating stage. In the future, this process could be seamlessly integrated into conventional CVD reactors by equipping them with an electron gun featuring programmable position control.
@@
Unlike many other area selective CVD processes, direct electron beam writing for seeding is not affected by the surface material or topography. To demonstrate this capability, the same precursor was used to fabricate vertical nanostructures with high aspect ratios using the focused electron beam. In addition to achieving autocatalytically active seeds at high aspect ratio structures, the first quantitative measurement of the temperature change during direct electron beam writing of such structures was performed. This quantitative determination of the up to now inaccessible local temperature evolution is a major breakthrough as it enables improved modelling for accurate shape 3D nanoprinting through FEBID within the additive manufacturing field at a sub-100 nm scale [46, 47, 48, 49, 50].
@@
To summarise, this work presents a new, technologically powerful and easy to implement approach to area selective gold deposition. The low temperature necessary makes it highly suitable for processing innovative flexible devices that rely on heat-sensitive polymers, including flexible electronics [52] and flexible solar cells [53]. Another exciting area of application is utilizing the maskless direct write seeding on 3D architectures, which enables the fabrication of nanostructures or localized electric contacts on intricate 3D devices.
4 Experimental Section
Silicon wafers with native oxide and different doping levels were employed as the substrates and chipped into 1 cm x 1 cm pieces. As the doping level decides the conductivity and therefore the secondary electron yields, the minimum required electron doses for successful seeding varied. The silicon chips with a native oxide layer were vertically sonicated in N-methyl-2-pyrrolidone for 20 min, followed by a rinse with isopropyl alcohol. The chips were then soaked in a 3:1 Piranha solution for at least 1 h before being rinsed with deionized water and dried using nitrogen gas. This process effectively removes any organic material from the surface of the sample.
All testing was carried out using a Helios 5 UX dual beam microscope. The gas injection system (GIS) with crystals of Au(acac)Me2 (CAS: 14951-50-9) was heated to 30 °C. To enable autocatalytic growth of gold films, gold nuclei were seeded with an electron beam at 5 kV acceleration voltage and a beam current of 100 pA. The base chamber pressure prior to opening the GIS valve was typically mbar, with the GIS valve open during autocatalysis mbar at the start of the experiments, dropping to approximately at the end. A Kleindiek vacuum compatible Micro Heating and Cooling Stage (MHCS) was used together with a temperature control system to heat the sample up to 110 °C. Pillars for measuring the temperature induced by an electron beam were grown using a focused electron beam of 15 kV and 100 pA. Deposition for all pillars was achieved by repeating a single spot 5725 times using the maximum possible dwell time of 26.2 ms, resulting in a total deposition time of about 2 min 30 s for a single pillar. The experimental matrices are fully described in the SI.
Lamellae for transmission electron microscopy (TEM) were prepared according to the standard lift-out protocol using the dual beam FIB-SEM system Tescan Lyra3. The TEM images were captured using a probe-corrected ThermoFisher Scientific Titan Themis 200 G3.
Supporting Information is available from the Wiley Online Library or from the author.
@@ Acknowledgements
This work was funded by the German Research Association under grant agreement no. HO 5461/3-1 and by the Swiss National Fund by the COST-SNF project IZCOZ0_205450. We want to furthermore acknowledge support by the EU COST action CA 19140 ’FIT4NANO’ (www-fit4nano.eu). The authors are particularly grateful to Caroline Hain for her support with the English language, and to Ibukun Olaniyan for the support with AFM measurements.
@@
1 Methods and Full Seed Array
Silicon wafers with native oxide and different doping levels were employed as the substrates and chipped into 1 cm x 1 cm pieces. As the doping level decides the conductivity and therefore the secondary electron yields, the minimum required electron doses for successful seeding varied. The silicon chips with a native oxide layer were vertically sonicated in N-methyl-2-pyrrolidone for 20 min, followed by a rinse with isopropyl alcohol. The chips were then soaked in a 3:1 Piranha solution for at least 1 h before being rinsed with deionized water and dried using nitrogen gas. This process effectively removes any organic material from the surface of the sample.
All testing was carried out using a Helios 5 UX dual beam microscope. The gas injection system (GIS) with crystals of Au(acac)Me2 (CAS: 14951-50-9) was heated to 30 °C. To enable autocatalytic growth of gold films, gold nuclei were seeded with an electron beam at 5 kV acceleration voltage and a beam current of 100 pA. The base chamber pressure prior to opening the GIS valve was typically mbar, with the GIS valve open during autocatalysis mbar at the start of the experiments, dropping to approximately at the end of an experimental day. A Kleindiek vacuum compatible Micro Heating and Cooling Stage (MHCS) was used together with a temperature control system to heat the sample up to 125 °C.
The complete pattern array for all experiments is shown in Figure SI.5(a). It comprises 2 µm x 2 µm squares with variable pixel-to-pixel pitches (5, 10 and 20 nm) of the electron beam, separated by 2 µm, and different deposition times (achieved by scan repeats from 1 to 100). A non-linear color scale indicates the deposition time of a particular square, ranging from 10 ms to 16 s. The whole seed array takes 1 min 18 s of deposition time to complete. The assortment of squares within the full seed array enables the gold film thickness to be optimised and to understand how autocatalytic growth depends on the electron beam patterning. Squares featuring the same deposition time (and dose) are highlighted with a coloured frame.
The experimental layout of all autocatalytically grown films is shown schematically in Figure SI.5(b). The complete experimental procedure for seeding and autocatalytic growth at various temperatures is detailed below:
@@ 1 Methods and Full Seed Array
Move the stage to the first patterning area, which should be at least 50 µm away from the focusing area (position i in Figure SI.5(b)), without conducting any observation with the SEM.
@@
Moving to the second patterning area, which is situated 100 µm from the previously grown films (position ii in Figure SI.5(b)).
@@
Pillars for measuring the temperature induced by an electron beam were grown using a focused electron beam of 15 kV and 100 pA. Deposition for all pillars was achieved by repeating a single spot 5725 times the maximum possible dwell time of 26.2 ms, resulting in a total deposition time of about 2 min 30 s for a single pillar. The layout for experimental deposition of all pillars under different stage temperatures is illustrated in Figure SI.5(c). The entire experimental protocol for pillars growth is described below:
@@
Without observation with the SEM, move the stage to the first patterning area, which is at least 20 µm away from the focusing area. Room temperature deposition is done at position i in Figure SI.5(c).
@@
Repeat the same protocol for higher temperatures ranging from 40 to 110 °C with in steps of 10 K.
@@
2 SE Micrographs of Autocatalytically Grown Films on the Full Seed Arrays
@@
SE micrographs of gold films, grown on the full seed arrays at various temperatures, are displayed in Figure SI.6. The gold films in the top row were grown on seed arrays deposited at a room temperature of 30 °C, while the seed arrays for the gold films in the bottom row were deposited at higher temperatures ranging from 60 to 90 °C. Autocatalytic growth was continued for 30 min for each film.
High-resolution SE image analysis of a gold film seeded and grown at 90 °C (Fig. SI.7) shows a lateral average grain size of 18 nm. Results are obtained with particle analysis plugin in open-source ImageJ software.
3 Time Evolution of Autocatalytic Growth on the Full Seed Arrays
@@
The autocatalytic growth of gold films on two seed arrays for additional intermediate electron supply is shown in Figure SI.8, indicating the time evolution. Both autocatalytic growth of gold on the seeds and electron beam irradiation processes are simultaneously involved in this process (similar to [12]). The initial SEM image at 0 minutes in the top left corner shows the as-deposited seed array at room temperature (30 °C), taken immediately after heating to 90 °C. With the GIS open and stage heated for 10 min, several images of this array were captured every 60 seconds. The images were taken with a field of view of 40 µm x 26.6 µm and a dwell time 15 µs. Thus, 100 pA electron beam current resulted in a dose of 0.022 pC/µm2 or 0.14 electrons per nm2 per single image. Figure SI.8 displays SEM images captured at 3, 7, and 10 mins to visualize the evolution of the autocatalytic gold CVD film growth. The bottom row presents the as-deposited seed array, which was grown on the sample at an elevated temperature of 90 °C, with the same sequence of images taken at different times. These images show that SE imaging acts as iterative seeding to enhance the CVD growth rate significantly compared to the pure autocatalytic growth demonstrated above. Note that when imaging, seeds are also deposited between squares, resulting in a decrease in area-selectivity. By implementing an iterative seeding approach, electrons will only be supplied in the selected areas.
4 Control Experiments